# for Large Language Models

### 06 Feb 2024 v2

@@ Abstract
In recent years, instruction tuning has gained increasing attention and emerged as a crucial technique to enhance the capabilities of Large Language Models (LLMs). To construct high-quality instruction datasets, many instruction processing approaches have been proposed, aiming to achieve a delicate balance between data quantity and data quality. Nevertheless, due to inconsistencies that persist among various instruction processing methods, there is no standard open-source instruction processing implementation framework available for the community, which hinders practitioners from further developing and advancing. To facilitate instruction processing research and development, we present [Uncaptioned image], an easy-to-use instruction processing framework for LLMs, which modularizes instruction generation, selection, and prompting, while also considering their combination and interaction. EasyInstruct is publicly released and actively maintained at https://github.com/zjunlp/EasyInstruct, along with a running demo App [^1] https://huggingface.co/spaces/zjunlp/EasyInstruct. for quick-start, calling for broader research centered on instruction data [^2] This is a subprobject of KnowLM (https://github.com/zjunlp/KnowLM), which facilitates knowledgeable LLM Framework with EasyInstruct, EasyEdit Wang et al. (2023a); Yao et al. (2023); Zhang et al. (2024), EasyDetect etc..
@@ 1 Introduction
To address this issue, we develop EasyInstruct as depicted in Figure 1, an easy-to-use instruction processing framework for LLMs. Given some existing chat data, corpus, or Knowledge Graphs, EasyInstruct can handle instruction generation, selection, and prompting processes, while also considering their combination and interaction. These consistencies facilitate further development and comparisons of various methods, thus promoting the advancement of better instruction processing work. We further conduct experiments with EasyInstruct to validate its effectiveness in instruction processing. We are committed to the long-term maintenance of EasyInstruct, providing continuous support for new features to ensure its effectiveness as a framework for instruction processing and synthetic data generation (Bauer et al., 2024).
@@ 3.1 Design Principles
The framework is designed to cater to users with varying levels of expertise, providing a user-friendly experience ranging from code-free execution to low-code customization and advanced customization options:
@@ 3.2 APIs & Engines
The score using the abstract syntax tree to encode structural and logical attributes, to evaluate the correlation between code and reasoning abilities. (Bi et al., 2023).
@@
Table 1: Components of Generators and Selectors modules of [Uncaptioned image]. The instruction generation methods implemented in Generators are categorized into three groups, based on their respective seed data sources: chat data, corpus, and knowledge graphs. The evaluation metrics in Selecors are divided into two categories, based on the principle of their implementation: statistics-based and LM-based.
@@ 3.3 Generators
Figure 2: A running example of instruction generation and selection in Refer to caption
Figure 2: A running example of instruction generation and selection in Refer to caption
@@ 5 Conclusion and Future Work
We present [Uncaptioned image], an easy-to-use instruction processing framework for LLMs. EasyInstruct can combine chat data, corpus, KGs and LLMs as an automated instruction generation tool, reducing the cost of manual data annotation. Additionally, EasyInstruct integrates a diverse set of instruction selection tools to optimize the diversity and distribution of instruction data, thereby improving the quality of fine-tuning data. EasyInstruct is designed to be easy to extend, and we will continue to update new features (e.g., knowledgeable synthetic data generation) to keep pace with the latest research. We expect EasyInstruct to be a helpful framework widely used by researchers and practitioners to facilitate their work of instruction tuning on LLMs.
6 Acknowledgements
We thank the developers of the self-instruct [^8] https://github.com/yizhongw/self-instruct library for their significant contributions to the NLP community. We thank the LLaMA team for providing us access to the models, and open-source projects, including Alpaca-LoRA [^9] https://github.com/tloen/alpaca-lora and AlpacaEval [^10] https://github.com/tatsu-lab/alpaca_eval.
@@ A.2.1 Shell Script
generator:
SelfInstructGenerator:
target_dir: data/generations/
data_format: alpaca
seed_tasks_path: data/seed_tasks.jsonl
generated_instructions_path: generated_instructions.jsonl
generated_instances_path: generated_instances.jsonl
num_instructions_to_generate: 100
engine: gpt-3.5-turbo
num_prompt_instructions: 8
@@
GPTScoreSelector:
%**** appendix.tex Line 125 ****engine: gpt-3.5-turbo

### 21 Mar 2024 v3

@@ Abstract
In recent years, instruction tuning has gained increasing attention and emerged as a crucial technique to enhance the capabilities of Large Language Models (LLMs). To construct high-quality instruction datasets, many instruction processing approaches have been proposed, aiming to achieve a delicate balance between data quantity and data quality. Nevertheless, due to inconsistencies that persist among various instruction processing methods, there is no standard open-source instruction processing implementation framework available for the community, which hinders practitioners from further developing and advancing. To facilitate instruction processing research and development, we present [Uncaptioned image] [^1] This is a subprobject of KnowLM (https://github.com/zjunlp/KnowLM), which facilitates knowledgeable LLM Framework with EasyInstruct, EasyEdit Wang et al. (2023a); Yao et al. (2023); Zhang et al. (2024), EasyDetect etc., an easy-to-use instruction processing framework for LLMs, which modularizes instruction generation, selection, and prompting, while also considering their combination and interaction. EasyInstruct is publicly released and actively maintained at https://github.com/zjunlp/EasyInstruct, along with an online demo app [^2] https://huggingface.co/spaces/zjunlp/EasyInstruct and a demo video [^3] https://youtu.be/rfQOWYfziFo for quick-start, calling for broader research centered on instruction data and synthetic data.
@@ 1 Introduction
To address this issue, we develop EasyInstruct as depicted in Figure 1, an easy-to-use instruction processing framework for LLMs. Given some existing chat data, corpus, or Knowledge Graphs, EasyInstruct can handle instruction generation, selection, and prompting processes, while also considering their combination and interaction. These consistencies facilitate further development and comparisons of various methods, thus promoting the advancement of better instruction processing work. We further conduct experiments with EasyInstruct to validate its effectiveness in instruction processing. Currently, EasyInstuct is open-sourced on GitHub and has already received around 300 stars. We are committed to the long-term maintenance of EasyInstruct, providing continuous support for new features to ensure its effectiveness as a framework for instruction processing and synthetic data generation (Bauer et al., 2024).
@@ 3.1 Design Principles
The framework is designed to cater to users with varying levels of expertise, providing a user-friendly experience ranging from code-free execution to low-code customization and advanced components extension options:
@@ 3.2 APIs & Engines
The score using the abstract syntax tree to encode structural and logical attributes, to evaluate the correlation between code and reasoning abilities (Bi et al., 2023).
@@
Table 1: Components of Generators and Selectors modules of [Uncaptioned image]. The instruction generation methods implemented in Generators are categorized into three groups, based on their respective seed data sources: chat data, corpus, and knowledge graphs. The evaluation metrics in Selecors are divided into four categories, based on the principle of their implementation: statistics-based, n-gram-based, structure-based, and LM-based.
@@ 3.3 Generators
from easyinstruct import SelfInstructGenerator
from easyinstruct import GPTScoreSelector
from easyinstruct.utils.api import set_openai_key
[…]
set_openai_key("YOUR-KEY")
[…]
generator = SelfInstructGenerator(
data_format = "alpaca",
seed_tasks_path = "seed_tasks.jsonl",
generated_instances_path = "generation.jsonl",
num_instructions_to_generate=100,
engine = "gpt-3.5-turbo",
[…]
generator.generate()
[…]
selector = GPTScoreSelector(
%**** design.tex Line 150 ****source_file_path = "generation.jsonl",
engine = "gpt-3.5-turbo",
threshold = 4,
[…]
selector.process()
Figure 2: A running example of instruction generation and selection in Refer to caption.
@@ 5 Conclusion and Future Work
We present [Uncaptioned image], an easy-to-use instruction processing framework for LLMs. EasyInstruct can combine chat data, corpus, KGs and LLMs as an automated instruction generation tool, reducing the cost of manual data annotation. Additionally, EasyInstruct integrates a diverse set of instruction selection tools to optimize the diversity and distribution of instruction data, thereby improving the quality of fine-tuning data. EasyInstruct is designed to be easy to extend, and we will continue to update new features (e.g., knowledgeable synthetic data generation) to keep pace with the latest research. We expect EasyInstruct to be a helpful framework for researchers and practitioners to facilitate their work of instruction tuning on LLMs.
@@ A.2.1 Shell Script
generator: SelfInstructGenerator: target_dir: data/generations/ data_format: alpaca seed_tasks_path: data/seed_tasks.jsonl generated_instructions_path: generated_instructions.jsonl generated_instances_path: generated_instances.jsonl num_instructions_to_generate: 100 engine: gpt-3.5-turbo num_prompt_instructions: 8
@@
%**** appendix.tex Line 125 ****GPTScoreSelector:
engine: gpt-3.5-turbo
@@ A.7 Example features in the Prompts module
A.8 Acknowledgements
We thank the developers of the self-instruct [^10] https://github.com/yizhongw/self-instruct library for their significant contributions to the NLP community. We thank the LLaMA team for providing us access to the models, and open-source projects, including Alpaca [^11] https://github.com/tatsu-lab/stanford_alpaca, Alpaca-LoRA [^12] https://github.com/tloen/alpaca-lora and AlpacaEval [^13] https://github.com/tatsu-lab/alpaca_eval.